OpenStudy (anonymous):

Can someone please help me with expressing a Complex Number in Polar Form? Without using a calculator? I'm having troubles with the unit circle. Thank you.

OpenStudy (anonymous):

The equation is z = -1 + 1i

OpenStudy (anonymous):

i found the modulus = r = | z | = sqrt(2); and I know that cos theta = -sqrt(2) / 2; but after that I am lost

OpenStudy (anonymous):

okay, and what do I do with that? I understand it as: :